OpenStudy (anonymous):

please help!! e^ln(6x)= a. ln(6) b. 6x c. ln(x) or d. 6ln(x) ??

OpenStudy (anonymous):

i think it is 6x but im not positive

OpenStudy (anonymous):

\[e^{\ln(\text{whatever})}=\text{whatever}\]

OpenStudy (anonymous):

so yes, it is \(6x\)

OpenStudy (anonymous):

thank you so much!!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

can you please help me with this one? im so confused here

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

log(10^3x^7)= a. 21 b. 3x^7 3. 21log(x) or d. 3+7log(x)

OpenStudy (anonymous):

\[\log(10^3x^7)\] like that?

OpenStudy (anonymous):

i got log(1000x^7) but that is not a choice

OpenStudy (anonymous):

yes just like th

OpenStudy (anonymous):

sorry i dont know how to do the math symbols on the computer haha

OpenStudy (anonymous):

no, your job is it split this up as \[\log(10^3)+\log(x^7)\] using \[\log(ab)=\log(a)+\log(b)\]

OpenStudy (anonymous):

then note that \[\log(10^3)=3\]

OpenStudy (anonymous):

and finally that \[\log(x^7)=7\log(x)\]

OpenStudy (anonymous):

making the whole thing \[3+7\log(x)\] which i hope is one of the choices

OpenStudy (anonymous):

ohh i understand now :)

OpenStudy (anonymous):

it is a choice! thank you so much i really appreciate it.. i have a huge test tomorrow and i couldnt ge

OpenStudy (anonymous):

i needed this for the study guide haha

OpenStudy (anonymous):

are you good with graphing? @satellite73

OpenStudy (anonymous):

the choices for this are a. ln(x-3)+4 b. ln(x)+4 c. ln(x+3) and d. ln(x+3)-4

OpenStudy (anonymous):

i think the answer

OpenStudy (anonymous):

i just need confirmation haha @satellite73

OpenStudy (anonymous):

whoops i didn't finish i think the answer is c. ln(x+3)